Uxx+3Uxy-4Uyy+Ux+4Uy=0 find the canonical form and solve the equation. (general solution)


1
Expert's answer
2021-02-01T16:47:14-0500

The characteristic quadratic form of this equation has the form

"Q( \lambda_1 , \lambda_2)=\lambda^2_1 +3 \lambda_1 \lambda_2-4 \lambda_2^2"

Let's bring it to its canonical form:

"\lambda^2_1 +3 \lambda_1 \lambda_2-4 \lambda_2^2=(\lambda_1+3/2 \lambda_2)^2-( \sqrt{11/2} \lambda_2)^2= k_1^2-k_2^2"

Where

"k1= \lambda_1+3/2 \lambda_2"

"k_2= \sqrt{11/2} \lambda_2"

So

"\begin{pmatrix}
 k_1 \\
 k_2 
\end{pmatrix} =\begin{pmatrix}
 1 & 3/2\\
 0 & \sqrt{11/2}
\end{pmatrix}\begin{pmatrix}
 \lambda_1 \\
 \lambda_2
\end{pmatrix}"

Find the number replacement matrix

"( \begin{pmatrix}
 1 & 3/2\\
 0 & \sqrt{11/2}
\end{pmatrix}^T)^{-1}=\begin{pmatrix}
 1 & 0 \\
 -30/47 & 20/47
\end{pmatrix}"


We replace the number:

"\begin{pmatrix}
 a \\
 b
\end{pmatrix}=\begin{pmatrix}
 1 & 0 \\
 -30/47 & 20/47
\end{pmatrix}\begin{pmatrix}
 x\\
 y
\end{pmatrix}"

a=x

b=(-30x+20y)/47

To substitute new variables in the original equation, we put


V(a,b)=u(x,y)

ux=va-(30/47)vb

uy=20/47 vb

uxx=vaa-(60/47)vab+900/2209vbb

uxy=(60/47)vab-1800/2209vbb

uyy=400/2209vbb


vaa-(60/47)vab+900/2209vbb+(180/47)vab-5400/2209vbb -1600/2209vbb +va-(30/47)vb+800/47 vb =vaa+(120/47)vab-6100/2209vbb + va+770/47vb=0


Answer: vaa+(120/47)vab-6100/2209vbb + va+770/47vb=0

hyperbolic type


General solution:

V(a,b)=f1(a)+f2(b)

u(x,y)=f1(x)+f2((-30x+20y)/47)